OpenStudy (anonymous):

Please heeeelp! a+b+c+d+e=200 a^2+b^2+c^2+d^2+e^2=1870 Please find the value of e. I have no idea how to solve these equations. The actual problem is different, but I have narrowed it down to this.

OpenStudy (anonymous):

@dmezzullo

OpenStudy (anonymous):

i think this is an inequality and the problem wants u to find possible values of e

OpenStudy (anonymous):

Hint: let \(x\) be a real number\[(a-x)^2+(b-x)^2+(c-x)^2+(d-x)^2+(e-x)^2=\]\[\ \ \ \ \ a^2+b^2+c^2+d^2+e^2-2(a+b+c+d+e)x+5x^2=1870-400x+5x^2\]

OpenStudy (anonymous):

@Juliette32801 @robtobey @satellite73 I reeally need help!

OpenStudy (anonymous):

Ill try to help you.. Wait one sec and ill be on my way (:

OpenStudy (anonymous):

Im sorry I dont know.. maybe some other question Good Luck (:

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

@amistre64 @dumbcow @kropot72

OpenStudy (anonymous):

Hmm. The only suggestion I could make is to solve each equation for e, then set them equal to each other. For example, from the first equation: e=200-a-b-c-d

OpenStudy (amistre64):

how do we know its been narrowed down correctly?

OpenStudy (anonymous):

Well, the narrowing down was just normal division, which I obviously know

OpenStudy (amistre64):

you have 2 equations in 5 unknowns; i see no way to pinpoint what "e" should be with this information

OpenStudy (amistre64):

if the question is asking you to solve for e given some original information, your method at a solution may not be well suited to the task. But without knowing the original information ... we really cant ascertain that.

OpenStudy (jasmineflvs):

Actually, these equations have no solution in the real numbers. Note that the arithmetic mean of the five numbers is 2005=40 while the quadratic mean is 18705−−−−√=19.339…, contradicting the QM−AM inequality.